# SPECIAL RELATIVITY

### v9

@@
Evgeny Saldin Draft from May 2024
@@
Evgeny Saldin Draft from May 2024
New material has been added to the Chapter 1, Chapter 3, Chapter 4, Chapter 5, Chapter 9, Chapter 10, Chapter 11, Chapter 13.
@@
There are many books on the classical subject of special relativity. However, after having spent a number of years, both in relativistic engineering and research with relativity, I have come to the conclusion that there exist a place for a new book. I do believe that the present book is not quite the same as the others, mainly due to attempt to cast light on dark corners.
I should make it clear what this little book is not. It is not a textbook on relativity theory. What the book is about is the nature of special relativistic kinematics, its relation to space and time, and the operational interpretation of coordinate transformations. Every theory contains a number of quantities that can be measured by experiment and an expressions that cannot possibly be observed. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. The distinction is not always made and many authors claim some data to be observable, according to arbitrary conventions, which do not correspond to any physical experiment. This leads to inconsistencies and paradoxes that should be avoided at all cost.
The practical approach used in the book should be acceptable to astronomers, space engineers, accelerator engineers, and more generally, relativistic engineers. This approach, unusual in the relativistic literature, may be clarified by quoting one of the problems discussed in the text: the new light beam kinematics for rotating frame of references. Since we live on such a rotating (earth-based) frame of reference, difference in relativistic kinematics between rotating and non-rotating frames of reference is of great practical as well as theoretical significance. A correct solution of this problem requires the use of relativistic principles even at low velocities since the first-order terms in $(v/c)$ play a fundamental role in the non-inertial relativistic kinematics of light propagation.
All the results presented here are derived from the ”first principles”, and all steps involving physical principles are given. To preserve a self-consistent style, I place the derivation of auxiliary results in appendices. To help readers form their own opinion on the topics discussed, the end of each chapter has a suggested bibliography together with relevant remarks. The list of references includes only the papers I have consulted directly. A lot of papers remain unmentioned, and for this I apologize.
@@ 1 Introduction
The standard books on special relativity do not usually address the questions of the physical meaning of relativistic effects and the nature of space-time. The presentation of the subject in the present book differs somewhat from the usual one in that the four-dimensional geometric formulation of the theory plays a dominant role than in most of the current textbooks.
The book begins with a critical survey of the present approaches to special relativity. The established way of looking at special relativity is based on Einstein postulates: the principle of relativity and the constancy of the velocity of light. In the most general geometric approach to the theory of special relativity, the principle of relativity, in contrast to Einstein formulation, is only a consequence of the geometry of space-time. We point that the essence of the special theory of relativity consists in the following postulate: all physical processes proceed in four-dimensional space-time, the geometry of which is pseudo-Euclidean.
@@
It should be emphasized that in practical applications there are two choices of clock synchronization convention useful to consider:
@@
There is a widespread view that only philosophers of physics discuss the issue of distant clock synchronization. Indeed, a typical physical laboratory contains no space-time grid. It should be clear that a rule-clock structure exist only in our mind and manipulations with non existing clocks in the special relativity are an indispensable prerequisite for the application of dynamics and electrodynamics theory in the coordinate representation. Such situation usually forces physicist to believe that the application of the theory of relativity to the study of physical processes is possible without detailed knowledge of the clocks synchronization procedure.
However, many problems in special relativity can be adequately treated only by an approach which uses the non-standard absolute time synchronization. One of the features that is unique to this book is its treatment of the absolute time coordinatization. No other books deals with Galilean transformations in the framework of special relativity.
@@
Third chapter presents an ”operational interpretation” of the Lorentz and absolute time coordinatizations. This is probably the most important chapter of this book. Today the statement about correctness of Galilean transformations is a ”shocking heresy”, which offends the ”relativistic” intuition and the generally accepted way of looking at special relativity of most physicists. The difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view will be an important discovery for every special relativity expert. To our knowledge, neither operational interpretation of the absolute time coordinatization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
We have tried to keep the mathematical complexity of the discussion to a minimum. The derivation of relativistic kinematics in the first part of the book (Chapter 3-11) is fairly elementary from a mathematical point of view, but it is conceptually subtle. We use synthetic approach to present the material: some simple models are studied first, and more complicated ones are introduced gradually.
@@
We start with aberration of light phenomena. Light, being a special case of electromagnetic waves, is described by the electrodynamics theory. It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a typical relativistic object as light.
@@
In the Chapter 4-5, we present a critical reexamination of the existing aberration of light theory. The phenomenon of aberration of light is by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the literature. The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we consider optical phenomena associated with a relative motion of two (or more) light sources. There is no problem to introduce Lorentz coordinates when these sources are independent. The possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate system (i.e. individual set of clocks) for each sources. Suppose now that the second (e.g. secondary) source interacts with the light beam radiated from the first source. The peculiarity of this problem with the viewpoint of the special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame.
@@
It is generally believed that for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved. This statement, presented in most textbooks, is incorrect. The first error in the textbook argument is conceptual in nature. The observer describes the light beam from the emitter at rest using Minkowski metric. According to textbooks, the metric of the moving mirror is the same (i.e. Minkowski metric). The authors of textbooks incorrectly belief that the common Lorentz coordinatization for emitter and mirror in the single inertial frame can be assigned. This is misconception. The second error in the textbooks follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of the tangential motion. In this case the ”mirror motion” is not a real observable effect. Indeed, if the mirror is infinite the problem is not time dependent. It is obvious that only the motion of a finite mirror is meaningful. All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e. frequency and wavenumber vector) but the group velocity. What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
@@
Suppose that we have a point source and observer with his measuring devices are at rest in the inertial lab frame. We demonstrate that when a finite aperture mirror is moving tangentially with a constant velocity and interacts in the far-field zone with the radiated light, there is a deviation of the energy transport for the reflected light. The result is extremely important from the applied point of view. According to the Babinet’s principle, this remarkable prediction of our aberration of light theory is correct also for light transmitted through a hole in the moving opaque screen or, consequently, through a moving open end of the telescope barrel.
@@
Questions related to transmission through the transversely moving pupil detection (e.g. a moving end of the telescope barrel) lead to serious misunderstanding, which is actually due to an inadequate understanding of several complicated aspects of the statistical optics. Today one is told that the phenomenon of aberration of light could be interpreted, using ray model of light. One case of rather great interest is that which corresponds to transmission through a moving telescope tube. The rays of light coming from the star are falling on the telescope tube and, according to literature, not interact with its sides. One could naively expect that the region of applicability of ray optics, following from the textbooks reasoning, should be identified with any spatially incoherent radiation. This is misconception. In particular, a spatially completely incoherent source (e.g. an incandescent lamp or a star) is actually a system of elementary (statistically independent) point sources with different offsets. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength. Radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individually incoherent sources. An elementary source produces in front of pupil detection effectively a plane wave and the measuring instrument always influences the measured radiation (due to unavoidable diffraction of a plane wave by an aperture). It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single point source characteristics like independence on source motion.
@@
The Chapter 6 deals with astronomical applications. The effect of stellar aberration seems to be one of the simplest phenomena in astronomical observations. In spite of its apparent simplicity, aberration seems to be one of the most intricate effects in special relativity. It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer. Observations show clearly that stellar aberration does not depend on the relative motion between star and telescope on the earth. The lack of symmetry, between the cases when either the source or telescope is moving is shown clearly on the basis of the separation of binary stars. The relative motion of these stars with respect to each other (and hence, with respect to the earth) is never followed by any aberration, although motion of these stars is, sometimes, much faster than that of the earth around the sun. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration. Contradiction is so plane that some physicists use this fact to argue that stellar aberration contradicts the special theory of relativity. There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions. We demonstrate that the fact that we do not see myriads of widely separated binaries in wild gyration does not require any fundamental change of outlook, but it does require that aberration of ”distant” stars is treated in the framework of space-time geometric approach.
@@
It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous of the observation of the oblique fall of raindrops by a moving observer. This is a classical kinematics method to the computation of the stellar aberration used in astronomy for about three hundred years. This book develops a relativistic theory of aberration of light in rotating frames. To see the consequences of relativistic modification of stellar aberration, we start with the classical aberration increment, or $\theta_{a}=v/V$. Here $v$ is the velocity of observer and $V$ is the velocity of particles. It is assumed that $v\ll V$. According to the conventional approach, if we neglect $v^{2}/c^{2}$ comparison with 1, stellar aberration increment is simplified to elementary formula $\theta_{a}=v/c$. This result is particularly remarkable because the study of stellar aberration is intimately connected with the old kinematics: the Galilean vectorial law of addition of velocities is actually used. We come to the conclusion that the standard analysis of the stellar aberration does not take into account the fundamental difference between velocity of light and velocity of raindrops.
@@
The satisfactory treatment of the relativistic modification should be based on the two relativistic parameters. For an observer on the earth, it is, with respect to the solar referential frame, of about 30 km/s, corresponding to the earth motion around the sun. Clearly, in the theory of stellar aberration, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$. In the theory of stellar aberration, we, however, must also consider the relativistic parameter $V/c$. For light $V/c=1$ and the special theory of relativity says that stellar aberration cannot be discussed as a topic in the classical theory. Light is always a (”ultra”) relativistic object, no matter how small the ratio $v/c$ may be. According to the special theory of relativity, physical processes takes place in a metric space-time geometry, so the geodesic interval must be used to describe aberration of light phenomenon.
@@
The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars. We demonstrate that for explanation of the aberration of light effect in the rotating frame of reference one does not need neither modify the special theory of relativity, nor apply general theory of relativity. It is only necessary to strictly follow the special relativity.
@@
All phenomena in non-inertial reference system should be considered in the framework of the space-time geometric approach using metric tensor. Using Langevin metric in a frame of reference attending rotation enables us to explain all earth-based experiments. A correct solution of this problem in the rotating frame requires the use of metric tensor even in experiments of first order in $v/c$ since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of a light beam propagation. The historical rate of the studies of the optical effects in rotating frames is very unusual. The Sagnac effect was discovered in 1913 and described by Langevin in 1921. It is interesting that the Langevin metric has in the past hardly ever been applied in the stellar aberration theory.
@@
We demonstrate that the aberration of light is a complex phenomenon which must be branched out into a number of varieties according to their origin. These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (with respect to the fixed stars) of the observer or of the light source. Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement.
@@
The main facts which a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems (they exhibit aberrations not different from other stars). We present here a theory which accounts for all these, and in addition gives new results. All earth-based experiments can be explained on the basis of the effect of the measuring instrument (i.e. physical unavoidable influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
@@
In the Chapter 8 we analyze the potential for exploiting earth-based sources in order to confirm the predictions of the relativistic aberration of light theory. We discussed a simple scaling model for the stellar aberration. The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift exist even in the case when star moves with the same velocity as the earth. We obtained a condition for optical similarity between the aberration of light from a distant star (which is moving with the earth velocity) and from the earth-based incoherent source. The proposed method of measuring the angle of aberration involves the use of earth-based sources and have a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth rotation axis. The apparent position of the source image is thus always a little displaced in the direction of the earth motion around the sun at that moment, and hence describe a small elliptical figure during the annular revolution the earth around its axis. In principle, records could be taken over a period of one day.
@@
It is generally believed that the special relativity is reciprocal theory. The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. In fact, without looking at anything external to the earth-based accelerated frame, one could determine the speed of the earth with respect to the sun-based frame by means of aberration of earth-based incoherent source measurements. No one has ever done such experiment, but we know what would happen from the astronomical observations.
@@
(3) It is always possible to chose such variables, in which metric of the accelerated frame will be diagonal. This is a consequence of the (pseudo-Euclidean) geometry of space-time.
@@
(4) The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames. It implies the elimination of the privilege frame, so meaning that the notion of absolute motion does not have physical content at all.
@@
We discuss this subject in the Chapter 9. In the argument above, steps (2) and (3) are correct, but step (1), and its consequence (4), are wrong. Argument (3) says that all inertial frames are equivalent with regard to physical laws, not with regard to physical facts. At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames. Where does the asymmetry comes from? The electrodynamics equations are not supplies all information about the physics of the situation. To solve the electrodynamics equations it is necessary to determine the initial conditions. The time after diagonalization is readily obtained by introducing the time offset factor. This time shift has the effect of rotation of the plane of simultaneity. As a consequence of this, the plane wavefront rotates in the accelerated frame after metric diagonalization. Now the information about the direction of observer acceleration is recorded in the initial conditions (radiation wavefront orientation).
@@
One can conclude that not all is relative in relativity, because this theory also contains some features which are absolute. Many people would like to think that since in the Lorentz coordinatization the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation. They would think that dynamics, based on the physical fields, is actually hidden in the language of pseudo-Euclidean geometry. However, it is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames. A resolution of the asymmetry paradox identifies inertial (pseudo-gravitational) force within the accelerated system as the agency of asymmetry. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamical methods. Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process. See Chapter 9 for more ideas along this line.
@@
There are several point to be made about the above results. One interesting question is, why we are not discuss about the aberration of light radiated by a laser? In regard to light aberration one should differentiate between that from the incoherent source and that from the laser source. One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using laser as a earth-based light source. Let us stress that the aberration of light cannot be measured using laser. In fact, the electromagnetic wave travels in the laser resonator forward and back reflecting from the mirrors. It could be said that the asymmetry cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator. We would like to emphasize a very important difference between laser and incoherent source. The radiation of incoherent source clearly depends on an initial conditions and there is influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. In contrast, when we are dealing with optical resonator the (3-space) boundary conditions define the energy propagation direction.
@@
One finds many books which say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration, because according to the principle of relativity only relative motion has any physical meaning. It follows that any observed asymmetry would lead to a contradiction with the principle of relativity. This argument is wrong. We discuss the apparent conflict between aberration of light and the principle of relativity in the Chapter 12. It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena. Not only in Einstein writings, but in every textbook on special theory of relativity can be found the formulation of principle of special relativity as follows: the laws of physics are the same in all inertial frames. Principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.

### v10

@@
Evgeny Saldin Final draft
@@
Evgeny Saldin Final draft
@@
There are many books on the classical subject of special relativity. However, after having spent a number of years, both in relativistic engineering and research with relativity, I have come to the conclusion that there exists a place for a new book. I do believe that the present book is not quite the same as the others, mainly due to the attempt to cast light on dark corners.
I should make it clear what this little book is not. It is not a textbook on relativity theory. What the book is about is the nature of special relativistic kinematics, its relation to space and time, and the operational interpretation of coordinate transformations. Every theory contains a number of quantities that can be measured by experiment and expressions that cannot possibly be observed. Whenever we have a theory containing an arbitrary convention, we should examine what parts of the theory depend on the choice of that convention and what parts do not. The distinction is not always made and many authors claim some data to be observable, according to arbitrary conventions, which do not correspond to any physical experiment. This leads to inconsistencies and paradoxes that should be avoided at all costs.
The practical approach used in the book should be acceptable to astronomers, space engineers, accelerator engineers, and more generally, relativistic engineers. This approach, unusual in the relativistic literature, may be clarified by quoting one of the problems discussed in the text: the new light beam kinematics for rotating frames of references. Since we live on such a rotating (earth-based) frame of reference, the difference in relativistic kinematics between rotating and non-rotating frames of reference is of great practical as well as theoretical significance. A correct solution of this problem requires the use of relativistic principles even at low velocities since the first-order terms in $(v/c)$ play a fundamental role in the non-inertial relativistic kinematics of light propagation.
All the results presented here are derived from the ”first principles”, and all steps involving physical principles are given. To preserve a self-consistent style, I place the derivation of auxiliary results in appendices. To help readers form their own opinions on the topics discussed, the end of each chapter has a suggested bibliography together with relevant remarks. The list of references includes only the papers I have consulted directly. A lot of papers remain unmentioned, and for this, I apologize.
@@ 1 Introduction
The book begins with a critical survey of the present approaches to special relativity. The established way of looking at special relativity is based on Einstein’s postulates: the principle of relativity and the constancy of the velocity of light. In the most general geometric approach to the theory of special relativity, the principle of relativity, in contrast to Einstein’s formulation, is only a consequence of space-time geometry. We point out that the essence of the special theory of relativity consists in the following postulate: all physical processes proceed in four-dimensional space-time, the geometry of which is pseudo-Euclidean. The presentation of the subject in the present book differs somewhat from the usual one in that the four-dimensional geometric formulation of the theory plays a dominant role.
@@
It should be emphasized that in practical applications there are two choices of clock synchronization conventions useful to consider:
@@
There is a widespread view that only philosophers of physics discuss the issue of distant clock synchronization. Indeed, a typical physical laboratory contains no space-time grid. It should be clear that a rule-clock structure exists only in our mind and manipulations with nonexisting clocks in special relativity are an indispensable prerequisite for the application of dynamics and electrodynamics theory in the coordinate representation. Such a situation usually forces physicists to believe that applying the theory of relativity to the study of physical processes is possible without detailed knowledge of the clock synchronization procedure. However, many problems in special relativity can be adequately treated only by an approach that uses the non-standard absolute time synchronization. One of the features that is unique to this book is its treatment of the absolute time coordinatization (i.e. Galilean transformations) in the framework of special relativity.
@@
The third chapter presents an ”operational interpretation” of the Lorentz and absolute time coordinatizations. This is probably the most important chapter of this article. Today the statement about the correctness of Galilean transformations is a ”shocking heresy”, which offends the ”relativistic” intuition and the generally accepted way of looking at special relativity of most physicists. The difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view will be an important discovery for every special relativity expert. To our knowledge, neither the operational interpretation of absolute time synchronization nor the difference between absolute time synchronization and Einstein’s time synchronization from the operational point of view, are given elsewhere in the literature.
@@
We start with the aberration of light phenomena. Light, being a special case of electromagnetic waves, is described by the electrodynamics theory. It is well known that the electrodynamics theory meets all requirements of the theory of relativity and therefore must accurately describe the properties of such a typical relativistic object as light.
@@
In Chapters 4-5, we present a critical reexamination of the existing aberration of light theory. The phenomenon of aberration of light is by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the literature. The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we consider optical phenomena associated with a relative motion of two (or more) light sources. There is no problem with introducing Lorentz coordinates when these sources are independent. The possibility to introduce Lorentz coordinates in this situation consists of introducing an individual coordinate system (i.e. individual set of clocks) for each source. Suppose now that the second (e.g. secondary) source interacts with the light beam radiated from the first source. The peculiarity of this problem with the viewpoint of the special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame.
@@
It is generally believed that for a mirror moving tangentially to its surface the law of reflection which holds for the stationary mirror is preserved. This statement, presented in most textbooks, is incorrect. The first error in the textbook argument is conceptual in nature. The observer describes the light beam from the emitter at rest using the Minkowski metric. According to textbooks, the metric of the moving mirror is the same (i.e. Minkowski metric). The authors of textbooks incorrectly believe that the common Lorentz coordinatization for emitter and mirror in the single inertial frame can be assigned. This is a misconception. The second error in the textbooks follows from the fact that the concept of a plane wave and an infinite plane mirror is used in the case of tangential motion. In this case the ”mirror motion” is not a real observable effect. Indeed, if the mirror is infinite the problem is not time-dependent. Only the motion of a finite mirror is meaningful. All experimental methods for measuring the aberration increment operate with light signals, and hence do not measure the phase velocity (i.e. frequency and wavenumber vector) but the group velocity. What authors of textbooks generally overlooked is the fact that the energy transport problem is well-defined only if the source and mirror apertures have already been specified.
@@
Suppose that we have a point source and observer with his measuring devices at rest in the inertial lab frame. We demonstrate that when a finite aperture mirror is moving tangentially with a constant velocity and interacts in the far-field zone with the radiated light, there is a deviation of the energy transport for the reflected light. The result is extremely important from the applied point of view. According to Babinet’s principle, this remarkable prediction of our aberration of light theory is correct also for light transmitted through a hole in the moving opaque screen or, consequently, through a moving open end of the telescope barrel.
@@
Questions related to transmission through the transversely moving pupil detection (e.g. a moving end of the telescope barrel) lead to serious misunderstanding, which is due to an inadequate understanding of several complicated aspects of the statistical optics. Today one is told that the phenomenon of aberration of light could be interpreted, using a ray model of light. One case of rather great interest is that which corresponds to transmission through a moving telescope tube. The rays of light coming from the star are falling on the telescope tube and, according to the literature, do not interact with its sides. One could naively expect that the region of applicability of ray optics, following the textbook’s reasoning, should be identified with any spatially incoherent radiation. This is a misconception. In particular, a spatially completely incoherent source (e.g. an incandescent lamp or a star) is a system of elementary (statistically independent) point sources with different offsets. The order of magnitude of the dimensions of the elementary statistically independent source is about $\lambda$, where $\lambda$ is the (visible) radiation wavelength. A radiation field generated by a completely incoherent source can be seen as a linear superposition of fields of individually incoherent sources. An elementary source produces in front of pupil detection effectively a plane wave and the measuring instrument always influences the measured radiation (due to the unavoidable diffraction of a plane wave by an aperture). It should be remarked that any linear superposition of radiation fields from elementary point sources conserves single-point source characteristics like independence on source motion.
@@
Chapter 6 deals with astronomical applications. The effect of stellar aberration seems to be one of the simplest phenomena in astronomical observations. Despite its apparent simplicity, aberration seems to be one of the most intricate effects of special relativity. It is widely believed that stellar aberration depends on the relative velocity of the source (star) and observer. Observations show clearly that stellar aberration does not depend on the relative motion between stars and telescopes on the earth. The lack of symmetry, between the cases when either the source or telescope is moving is shown clearly on the basis of the separation of binary stars. The relative motion of these stars with respect to each other (and hence, with respect to the earth) is never followed by any aberration, although the motion of these stars is, sometimes, much faster than that of the earth around the sun. It should be stressed that it is the telescope and not the star that must change its velocity (relative to the fixed stars) to cause aberration. Contradiction is so plain that some physicists use this fact to argue that stellar aberration contradicts the special theory of relativity. There is no available explanation for the fact that, while the observational data on stellar aberration are compatible with moving earth, the symmetric description, when the star possesses the relative transverse motion, does not apparently lead to observations compatible with predictions. We demonstrate that the fact that we do not see myriads of widely separated binaries in wild gyration does not require any fundamental change of outlook, but it does require that the aberration of ”distant” stars is treated in the framework of space-time geometric approach.
@@
It is generally believed that the phenomenon of aberration of light could be interpreted, using the corpuscular model of light, as being analogous to the observation of the oblique fall of raindrops by a moving observer. This is a classical kinematics method for the computation of the stellar aberration used in astronomy for about three hundred years. This article develops a relativistic theory of aberration of light in rotating frames. To see the consequences of relativistic modification of stellar aberration, we start with the classical aberration increment $\theta_{a}=v/V$. Here $v$ is the velocity of the observer and $V$ is the velocity of particles. It is assumed that $v\ll V$. According to the conventional approach, if we neglect $v^{2}/c^{2}$ comparison with 1, stellar aberration increment is simplified to elementary formula $\theta_{a}=v/c$. This result is particularly remarkable because the study of stellar aberration is intimately connected with the old kinematics: the Galilean vectorial law of addition of velocities is used. We conclude that the standard analysis of the stellar aberration does not take into account the fundamental difference between the velocity of the light and the velocity of raindrops.
@@
The satisfactory treatment of the relativistic modification should be based on the two relativistic parameters. For an observer on the earth, it is, with respect to the solar referential frame, of about 30 km/s, corresponding to the earth’s motion around the sun. Clearly, in the theory of stellar aberration, we consider the small expansion parameter $v/c\simeq 10^{-4}$ neglecting terms of order of $v^{2}/c^{2}$. In the theory of stellar aberration, we, however, must also consider the relativistic parameter $V/c$. For light $V/c=1$ and the special theory of relativity says that stellar aberration cannot be discussed as a topic in the classical theory. Light is always a (”ultra”) relativistic object, no matter how small the ratio $v/c$ may be. According to the special theory of relativity, physical processes take place in a metric space-time geometry, so the geodesic interval must be used to describe the aberration of light phenomenon.
@@
The problem of the earth-based measurements is solved with the discovery of the essential asymmetry between the earth-based and the sun-based observers, namely, the acceleration of the traveling earth-based observer relative to the fixed stars. We demonstrate that for an explanation of the aberration of light effect in the rotating frame of reference, one does not need to modify the special theory of relativity nor apply the general theory of relativity. It is only necessary to strictly follow the special relativity.
@@
All phenomena in non-inertial reference systems should be considered in the framework of the space-time geometric approach using metric tensor. Using Langevin metric in a frame of reference attending rotation enables us to explain all earth-based experiments. A correct solution of this problem in the rotating frame requires the use of metric tensor even in experiments of the first order in $v/c$ since the crossed term in the Langevin metric (which involves the first-order deviation of metric tensor from its Minkowski form) plays a fundamental role in the non-inertial kinematics of light beam propagation. The historical rate of the studies of the optical effects in rotating frames is very unusual. The Sagnac effect was discovered in 1913 and described by Langevin in 1921. It is interesting that the Langevin metric has in the past hardly ever been applied in the stellar aberration theory.
@@
We demonstrate that the aberration of light is a complex phenomenon that must be branched out into a number of varieties according to their origin. These branching out takes place depending on what is the cause of aberration - whether it is a jump in the velocity (with respect to the fixed stars) of the observer or of the light source. Aberration could undergo further splitting - depending on the physical influence of the optical instrument on the measurement.
@@
The main facts that a theory of stellar aberration in the earth-based frame of reference must explain are (1) the annual apparent motion of the fixed stars about their locations and (2) the null apparent aberration of rotating binary systems (they exhibit aberrations not different from other stars). We present here a theory that accounts for all these, and in addition, gives new results. All earth-based experiments can be explained based on the effect of the measuring instrument (i.e. physical unavoidable influence of the telescope on the measurement) and the acceleration of traveling earth-based observer relative to the fixed stars.
@@
In Chapter 10 we analyze the potential for exploiting earth-based sources to confirm the predictions of the relativistic aberration of light theory. We discussed a simple scaling model for the stellar aberration. The motion of the stars with respect to the earth is never followed by any aberration. The aberration shift exists even in the case when a star moves with the same velocity as the Earth. We obtained a condition for optical similarity between the aberration of light from a distant star (which is moving with the earth’s velocity) and from the earth-based incoherent source. The proposed method of measuring the angle of aberration involves the use of earth-based sources and has a big advantage. The rotation of the earth on its axis should produce a corresponding shift of the image. The aberration shift depends only upon the value of $v_{\perp}$, the component of the orbital velocity perpendicular to the earth’s rotation axis. The apparent position of the source image is thus always a little displaced in the direction of the earth’s motion around the sun at that moment, and hence describes a small elliptical figure during the annular revolution of the earth around its axis. In principle, records could be taken over one day.
@@
It is generally believed that special relativity is a reciprocal theory. The aberration of light problem in the accelerated frame demonstrated the essential asymmetry between the accelerated and inertial observers. In fact, without looking at anything external to the earth-based accelerated frame, one could determine the speed of the earth with respect to the sun-based frame using aberration of earth-based incoherent source measurements. No one has ever done such an experiment, but we know what would happen from the astronomical observations.
@@
(3) It is always possible to choose such variables, in which the metric of the accelerated frame will be diagonal. This is a consequence of the (pseudo-Euclidean) geometry of space-time.
@@
(4) The validity of Maxwell’s laws in all inertial frames implies that light propagates with the same velocity $c$ in all inertial frames. It implies the elimination of the privileged frame, so meaning that the notion of absolute motion does not have physical content at all.
@@
In the argument above, steps (2) and (3) are correct, but steps (1), and its consequence (4), are wrong. Argument (3) says that all inertial frames are equivalent with regard to physical laws, not with regard to physical facts. At first glance, after the diagonalization of the Langevin metric, we have the symmetry between inertial frames. Where does the asymmetry come from? The electrodynamics equations do not supply all information about the physics of the situation. To solve the electrodynamics equations it is necessary to determine the initial conditions. The time after diagonalization is readily obtained by introducing the time offset factor. This time shift has the effect of rotation of the plane of simultaneity. As a consequence of this, the plane wavefront rotates in the accelerated frame after metric diagonalization. Now the information about the direction of observer acceleration is recorded in the initial conditions (radiation wavefront orientation).
@@
One can conclude that not all is relative in relativity because this theory also contains some features which are absolute. Many people would like to think that since in the Lorentz coordinatization, the time and distance have direct physical meaning, there should be some physical (dynamical) reason for this wavefront rotation. They would think that dynamics, based on the physical fields, is hidden in the language of pseudo-Euclidean geometry. However, it is a dynamical line of arguments that explains this paradoxical situation with the asymmetry between inertial and accelerated reference frames. A resolution of the asymmetry paradox identifies inertial (pseudo-gravitational) force within the accelerated system as the agency of asymmetry. The principle of equivalence can be applied to solve non-inertial kinematic problems with dynamical methods. Wavefront rotation associated with the transformation from the inertial frame to the accelerated (with respect to the fixed stars) frame may be regarded as a result of the action of pseudo-gravitational potential gradient during the acceleration process.
@@
There are several points to be made about the above results. One interesting question is, why we are not discussing the aberration of light radiated by a laser? Regarding light aberration, one should differentiate between that from the incoherent source and that from the laser source. One could naively expect that the rotation of the earth around the sun produces aberration in an amount large enough to be taken into account in precise observation work using a laser as an earth-based light source. Let us stress that the aberration of light cannot be measured using a laser. The electromagnetic wave travels in the laser resonator forward and back reflecting from the mirrors. It could be said that the asymmetry cancels during the (round-trip-to-round-trip) evolution of the radiation in the optical resonator. We would like to emphasize a very important difference between laser and incoherent source. The radiation of an incoherent source clearly depends on initial conditions and there is an influence of the mixture of positions and time which is a result of pseudo-gravity experienced by the accelerated observer. In contrast, when we are dealing with an optical resonator the (3-space) boundary conditions define the energy propagation direction.
@@
One finds many books that say that we cannot distinguish by any experiment which observer remains ”at rest” during the acceleration because according to the principle of relativity, only relative motion has any physical meaning. It follows that any observed asymmetry would lead to a contradiction with the principle of relativity. This argument is wrong. It is demonstrated that there is no conflict between the fundamental structure of special relativity on the one hand, and the aberration of light phenomena. Not only in Einstein’s writings, but in every textbook on the special theory of relativity can be found the formulation of the principle of special relativity as follows: the laws of physics are the same in all inertial frames. The principle of special relativity is irrelevancy of velocity with regard to physical laws, not with regard to anything.